2. 120 castings are produced by die casting process. The castings are inspected to check their surface roughness and dimensional accuracy. The results from 120 castings are summarized as the following: Dimensional Accuracy good 85 10 Bad 15 10 Surface Roughness High Low Let A denote the event that a casting has high surface roughness, and let B denote the event that a casting has bad dimensional accuracy. Determine the number of castings in a) An B b) An B c) AU B d) AUB
Answer #1

(a)

Dimensional Accuracy good (B') Dimensional accuracy bad (B) Total
Surface roughness high (A) 85 15 100
Surface roughness low(A') 10 10 20
Total 95 25 120

Aigcap B = Number of castings with high surface roughness AND bad dimensional accuracy = 15

So,

Answer is:

15

(b)

AB = Number of castings with high surface roughness AND good dimensional accuracy = 85

So,

Answer is:

85

(c)

Aigcup B = Number of castings with high surface roughness OR bad dimensional accuracy

=A+B-Aigcap B

= 100 + 25 - 15

= 110

So,

Answer is:

110

(d)

AB 4 = Number of castings with low surface roughness OR bad dimensional accuracy

  =A'+B-A'igcap B

= 20 + 25 - 10

= 35

So,

Answer is:

35